Strategic Initiatives and CSR

Get Hitched Inc. is a production company that is in the process of testing a strategic initiative aimed at increasing gross profit. The company’s current sales revenue is $1,500,000. Currently, the company’s gross profit is 35% of sales, but the company’s target gross profit percentage is 40%. The company’s current monthly cost of production is $975,000. Of this cost, 60% is for labor, 20% is for materials, and 20% is for overhead.

The strategic initiative being tested at Get Hitched is a redesign of its production process that splits the process into two sequential procedures. The make up of the costs of production for Procedure 1 is currently 50% direct labor, 45% direct materials, and 5% overhead. The makeup of the costs of production for Procedure 2 is currently 55% direct labor, 25% direct materials, and 20% overhead. Company management estimates that Procedure 1 costs twice as much as Procedure 2.

1. Determine what the cost of labor, materials, and overhead for both Procedures 1 and 2 would need to be for the company to meet its target gross profit at the current level of sales.

Cost makeup of Procedure 1:

Direct Labor $
Direct Materials
Overhead
Total $

Cost makeup of Procedure 2:

Direct Labor $
Direct Materials
Overhead
Total $

2. The company’s actual direct materials cost is $279,000 for Procedure 1. Determine the actual cost of direct labor, direct materials, and overhead for each procedure, and the total cost of production for each procedure.

Cost makeup of Procedure 1:

Direct Labor $
Direct Materials
Overhead
Total $

Cost makeup of Procedure 2:

Direct Labor $
Direct Materials
Overhead
Total $

3. The company is planning a CSR initiative to reuse some of the indirect materials used in production during Procedure 2. These indirect materials normally make up 60% of the overhead cost for Procedure 2, but the CSR initiative would reduce the usage of indirect materials. Determine what the maximum new cost of these indirect materials could be for Procedure 2 if this CSR initiative is expected to enable the company to meet its target gross profit percentage (holding all other costs constant).

Maximum new cost of P2 overhead materials:
$?
Answer #1
Answer
Working Note For-1
Current sales level $      15,00,000
Less: Target g/p(40%*sales) $        6,00,000
Total cost(P1+P2) $      12,60,000
Total cost P1(900000*2/3) $        6,00,000
Total cost P2 (900000*1/3) $        3,00,000
Answer-1 1…
Cost makeup of Procedure 1:
Direct Labor(600000*50%) $        3,00,000
Direct Materials(600000*45%) $        2,70,000
Overhead (600000*5%) $           30,000
Total $        6,00,000
Cost makeup of Procedure 2:
Direct Labor (300000*55%) $        1,65,000
Direct Materials(300000*25%) $           75,000
Overhead (300000*20%) $           60,000
Total $        3,00,000
Answer-2 2..
Cost makeup of Procedure 1:
Direct Labor(279000/45%*50%)= $        3,10,000
Direct Materials (Given) $        2,79,000
Overhead(279000/45%*5%)= $           31,000
Total $        6,20,000
Cost makeup of Procedure 2:
Direct Labor (310000*55%)= $        1,70,500
Direct Materials(310000*25%)= $           77,500
Overhead (310000*25%)= $           62,000
Total $        3,10,000
Answer-3 3..Forming an equation
Total cost for Both procedures+Gross profit=Current Sales
Assuming OH cost for procedure 2 as x,
620000+170500+77500+ x+600000=1500000
x=1500000-620000-170500-77500-600000
32000
ANSWER:
Maximum new cost of P2 overhead materials: $ 32,000
Please hit LIKE button if this helped.